Constants Part A Coulombs law for the magnitude of the torce F between twco parsicles with charges Q and Q separated by a distance d is What is the net orce exerted by these two charges on a thied charge 9s-50.5 nC placed betweeng Your answer may be positive or negative, depending on the direction of the force. Express your answer numerieally in newtons to three significant flgures. View Available Hints) and -1.130? where 8854 x 10-* c/(N m2) s the permittivity of tree space Consider two point charges located on the x axis: one charge, 1125 DC, is located at -1.710 m; the second charge, 9s-30.0 nC is at the origin (Z-0.0000) Fore on qs--32.106 Submit incorrect; Try Again; 11 attempts remaining
Answer #1

Given : q1 = -12.5 nC , r1 = -1.710 m , q2 = 38 nC , r2=0 m ,

q3 = 50.5 nC , r3 = -1.130 m

Constant : K = 9×109 N

Solution:

Part A Net force on charge q3 due to q1 and q2

As we know that , the force between two charges q and q' is given by:

F = K|qq'|/r2 , where r is the distance between two charges.

Distance between charge q1 and q3 is given by:

r = r1-r3 = -1.710-(-1.130) = 0.58 m

Force on charge q3 due to q1 is given by:

F13 = K|q1q3|/r2

= (9×109)(-12.5×10-9)(50.5×10-9)/(0.58)2

= 1.69 × 10-5 N ( it is direcring toward negative x axis )

Distance between charge q2 and q3 is 1.130 m

Force on charge q3 due to q2 is given by:

F23 = K(q2q3)/r2

=(9×109)(38×10-9)(50.5×10-9)/(1.130)2

= 1.35×10-5 N (it is directing toward negative x axis)

Net force acting on charge q3 :

Fnet = F13+F23

= (1.69×10-5 N)+(1.35×10-5 N)

= 3.04 × 10-5 N

As it is directed toward negative x axis so , its sign will be negative.

Hence Fnet = -3.04×10-5 N